OpenStudy (anonymous):

Lucy has 36 coins consisting of nickels and quarters amounting to $4.80. How many coins of each kind does she have?

OpenStudy (anonymous):

x+y=36 5x+25y=480 which, divided by 5, is x+5y=96. Subtract the first equation from teh second, and get 4y=60, so y=15. 15 quarters and 21 nickels. does that work? add to check

OpenStudy (anonymous):

yes, it works. just checked.

OpenStudy (anonymous):

okay thanks, i had 15 nickels & 21 quarters so it was the other way around